OpenStudy (anonymous):

PLEASE HELP ME!! Which of the following is not true based on the fact that 2/5x divided by 3x^2 = 8/15x^3? A. 3x^2/4 * 8/15x^3 = 2/5x B. 8/15x^3 * 3x^2/4 = 2/5x C. 2/5x divided by 8/15x^3 = 3x^2/4

OpenStudy (anonymous):

so you have a/b=c so cxb=a

OpenStudy (anonymous):

Can you figure it out now?

OpenStudy (anonymous):

Umm, no. Sorry, that didn't make any sense!

OpenStudy (anonymous):

do you understand the equation a/b=c so cxb=a?

OpenStudy (anonymous):

Basically to get a alone you have to multiply c by b

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

So now lets tage our variable pretend a=2/5x b=3x^2 and c=8/15x^3

OpenStudy (anonymous):

so a/b=c (2/5x)/(3x^2)=8/15x^3 a b c so now what is b times c? (8/15x^3) times (3x^2)= a and what is a? (2/5x)

OpenStudy (anonymous):

So the answer is?

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

Actually its B

OpenStudy (anonymous):

ohhhh, ok, I get what you're saying! Thanks!!

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Actually.. Wouldn't "B" be one of the correct answers?

OpenStudy (anonymous):

Yes that is what I mean B is the answer to your multiple choice problem

OpenStudy (anonymous):

the question is which of the following is NOT true... So you're saying B wouldn't work in this equation?

OpenStudy (anonymous):

I am so sorry I missed that part you were right the first time then. The answer is c c is not possible

OpenStudy (anonymous):

I totally didn't see the NOT part and just went on trying to explain how to understand the problem

OpenStudy (anonymous):

haha, that's okay!! I thought I was going insane for a minute there!